OpenStudy (anonymous):

Radians to degrees. Convert pi/6 into degrees. Please show work, I need to learn how to get the answer.

OpenStudy (anonymous):

pi is the equivalent of 180 degrees. so if you want it in degrees its basically 180/6 so 30 degrees

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (anonymous):

basically whenever you see pi think 180 degrees and whenever you see 180 degrees think pi

OpenStudy (anonymous):

k, I will, your tip made this a lot easier!

OpenStudy (anonymous):

no prob :)

OpenStudy (johnweldon1993):

\[\frac{ \pi }{ 6 } * \frac{ 180 }{ \pi }\] to convert from radians to degrees...you multiply by 180/pi to convert from degrees to radians....you multiply by pi/180 just another way of looking at it :)